OpenStudy (yumeybitaa):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21. −45 0 2 63

OpenStudy (johnweldon1993):

\[\large f(x) = 6(x + 2) - 3\] First we need to switch the 'x' and y' \[\large x = 6(y + 2) - 3\] Now we just solve for 'y' again \[\large x = 6y + 12 - 3\] \[\large x = 6y + 9\] \[\large 6y = x - 9\] \[\large y = \frac{x - 9}{6}\] That is our inverse \[\large f(x) = 6(x + 2) - 3\] \[\large f^{-1}(x) = \frac{x - 9}{6}\] Now just plug in 21 for 'x' and solve for 'y'

OpenStudy (anonymous):

i never understood the wording of these questions although i see them all the time here "solve for the inverse function when x = 21. "

OpenStudy (johnweldon1993):

^Indeed, the wording is terrible, but it always seems to be the same thing they are asking

OpenStudy (anonymous):

i can only assume it means find \[f^{-1}(21)\] which translates as "solve for \(x\) \[\large 6(x+2) − 3=21\]

OpenStudy (johnweldon1993):

Ha what do you know it comes out to the same answer

OpenStudy (anonymous):

add 3 divide by 6 subtract 2